OpenStudy (anonymous):

The equation below gives the height h of a passenger on a Ferris wheel at any time t during the ride to be h = 134 − 124 cos π/10t where h is given in feet and t is given in minutes. Use this equation to find the times at which a passenger will be 80 feet above the ground during the first revolution. Round your answers to the nearest tenth of a minute. Use your graphing calculator to graph the function and verify your answers. (Enter your answers as a comma-separated list.)

OpenStudy (anonymous):

80=134-124cos pi/10t -54=124cos (pi/10)t arccos(-54/124)=(pi/10)t Where do I go from there?